OpenStudy (anonymous):

find the x and y intercepts of 3x^(2/3)-2x. show all work please

OpenStudy (asnaseer):

\[y = 3x ^{2/3} - 2x\] y intercept is when x = 0, so y intercept is y = 0 x intercept is when y = 0 giving \[0 = 3x ^{2/3} - 2x = x(3x ^{-1/3} - 2)\] so x = 0 or \[3x ^{-1/3} - 2 = 0\] 2nd solution can be re-written as \[3x ^{-1/3} = 2\] \[3 = 2x ^{1/3}\] \[x ^{1/3} = 3/2\] \[x = (3/2) ^ {3} = 27/8 = 3.375\] so x intercepts are x = 0 and x = 3.375